1.Which ratio represent the number (PI, 3,14) ( One of them are correct)
A) Diameter:Radius
B) Diameter: Circumference
C) Circumference: Diameter
D) Circumference: Radius
THANK YOU IN ADVANCE

Answer by drk(1908)

Circumference = 2*pi*r
diameter = 2r
C/d = 2*pi*r/(2*r) = pi
Circumference / diameter is answer, [ C ]